OpenStudy (anonymous):

what are the holes, vertical asymptotes, and horizontal asymptotes of f(x)=x^3-7x^2+12x/ 3x^2-15x+12 ?

OpenStudy (anonymous):

Your VA will be where the denominator equals zero and your HA you simplify the highest powers. x^3/3x^2=x/3, which means you do not have a HA asymptote because it is constantly getting bigger

OpenStudy (anonymous):

so what exactly r the answers?

OpenStudy (anonymous):

VA=(3x-3)(x-4) VA=1 and 4

OpenStudy (anonymous):

Do you understand how to get the answer? If not I can explain it

OpenStudy (anonymous):

Yes i do now. Thanks so much!

OpenStudy (anonymous):

do u know how to solve.... \[\log _{16}3n=\log _{16}(2n+1)\] ?

OpenStudy (anonymous):

Yes, one sec and i will explain it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok so you will need to take the base 16 log of both sides, which will get rid of the logs 3n=2n+1 n=1

OpenStudy (anonymous):

I am not familiar with the equation drawer on here or I would use that. :) Does this make sense?

OpenStudy (anonymous):

Im referring to the problem

OpenStudy (anonymous):

yep!!

OpenStudy (anonymous):

 how do you know if this expression is "undefined" or not?